How can i solve this equation;
7x+5y=9
4x+5y=6

Answer by AAfter.Com(18)

7x+5y=9
4x+5y=6
--------
(subtracting) 3x = 3
=> x = 1
y = (9 - 7X1)/5 = 2/5.